OpenStudy (campayne):

Find the zeros of the function by using the Quadratic Formula. f(x) = x^2 + 3x - 7

OpenStudy (3mar):

If I may help?

OpenStudy (campayne):

PLEASE

OpenStudy (3mar):

Thanks!

OpenStudy (3mar):

Are you familiar with the Quadratic Formula?

OpenStudy (campayne):

not even a bit

OpenStudy (3mar):

Take that and tell me after doing math!

OpenStudy (3mar):

أنا ورايا أكل

OpenStudy (campayne):

this is still making zero sense

OpenStudy (eliesaab):

In \[ x^2 + 3x - 7=\\ ax^2 + bx + c =0 \] What is a,b, and c?

OpenStudy (eliesaab):

a=1; b=? c=?

OpenStudy (eliesaab):

Are you with me?

OpenStudy (campayne):

yes

OpenStudy (eliesaab):

What is b and c?

OpenStudy (campayne):

3 and 7?

OpenStudy (eliesaab):

3 is correct but 7 is not

OpenStudy (campayne):

um 1?

OpenStudy (eliesaab):

c=-7

OpenStudy (eliesaab):

Knowing a,b and c, you find the roots by using \[ x_1=\frac{-b -\sqrt{b^2-4 a c}}{2 a}\\ x_2=\frac{-b +\sqrt{b^2-4 a c}}{2 a} \]

OpenStudy (campayne):